Suppose the coupon rate for a TIPS is 2.6% and that you’ve just today purchased a new $10,000 face value issue. Assume that the annual inflation rate over the next year is 2.2%. a) After 6 months, what is the inflation adjusted principle of the TIPS? b) What is the dollar value of the coupon paid at the end of those first six months? c) At the end of the year, what are the inflation adjusted principle and the dollar value of the coupon payment?
Answer #1

Face Value (FV)= $10,000

Coupon rate (r)= 2.6%

Annual; inflation rate (i)=2.2%

(a) Inflation adjusted principal after 6 months= FV*(1+i/2)= $10,000*1+1.1%)= $10,000*1.011= $10,110

(b) Dollar value of coupon paid at the end of first 6 months= Inflation adjusted Principal*r/2 =$10,110*2.6%/2

=$10,110*0.013= $131.43

(c) Inflation adjusted principal at the end of one year= FV*(1+i )= $10,000*(1+2.2%)= $10,220

Dollar value of coupon paid at the end of one year= Inflation adjusted Principal*r/2 = $10,220*2.6%/2

=$10,220*0.013= $132.86